# CFG++: Manifold-constrained Classifier Free Guidance for Diffusion Models

### v1

@@ 3.2 Geometry of CFG++
Algorithm 3 DDIM Inversion with CFG
1:${\boldsymbol{x}}_{0},0\leq{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\omega}\in\mathbb{R}$
2:for $i=0$ to $T-1$ do
4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{x}}_{t+1}=\sqrt{\bar{\alpha}_{t+1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t+1}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$
6:end for
7:return ${\boldsymbol{x}}_{T}$
Algorithm 4 DDIM Inversion with CFG++
${\boldsymbol{x}}_{0},{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\lambda}\in[0,1]$
2:for $i=0$ to $T-1$ do
4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{x}}_{t+1}=\sqrt{\bar{\alpha}_{t+1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t+1}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$
6:end for
7:return ${\boldsymbol{x}}_{T}$
Figure 6: Comparison between DDIM inversion. CFG++ proposes a simple yet effective fix: using ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ in Tweedie’s denoising step.
@@
Figure 7: T2I using SD v1.5, CFG vs CFG++ ($\omega=9.0,\lambda=0.8$). Unnatural depictions of human hands, and incorrect renderings of the text by CFG are corrected in CFG++
@@
Table 1: Quantitative evaluation (FID, CLIP-similarity) of T2I with SD v1.5
@@ 4 Experimental Results
In this section, we design experiments to show the limitation of CFG and how CFG++ can effectively mitigate these downsides. We select Stable Diffusion (rombach2022high) (SD) as our baseline model for its widespread use. Experiments are conducted using SD v1.5 from diffusers (patil2022stable) unless specified otherwise. Notice that we use the conditioning scale $\lambda\in[0,1]$ for CFG++ whereas for standard CFG, the typical range is $\omega>1.0$. When searching for the matched guidance values of $\omega$ and $\lambda$ for a fair comparison, we fix $\lambda=0.2,0.4,0.6,0.8,1.0$ and find the $\omega$ values that produce the images that are of closest proximity in terms of LPIPS distance given the same seed. We found that the corresponding values were $\omega=2.0,5.0,7.5,9.0,12.5$, respectively, as shown in Appendix Fig. 14.
@@ 4.1 Text-to-Image Generation
Using the corresponding scales for $\omega$ and $\lambda$, we directly compare the performance of the T2I task using SD v1.5 and SDXL. In Tab. 1, we observe a constant improvement of the FID metric across all guidance scales, with approximately the same level of CLIP similarity or better. The improvements can also be clearly seen in Fig. 7 (SD v1.5), where the unnatural components of the generated images are corrected. Specifically, we see that unnatural depictions of human hands, and incorrect renderings of the text are corrected in CFG++, a long-standing research question in and of its own (podell2023sdxl; pelykh2024giving; chen2023textdiffuser).
Figure 8: Interpolating behavior of T2I along various guidance scales using SDXL with CFG++. The generated visuals align more significantly with the given textual description as the guidance scale increase to 1.
In Fig. 8, we visualize multiple images generated by CFG++ as we increase the guidance scale $\lambda$, illustrating a smooth transition from unconditional to highly conditional sampling. Specifically, at higher guidance scales, the target concept is more pronounced, with less emphasis on noisy, unrelated elements. This results in images that more closely reflect the specified text conditions. For example, in the second row, the images increasingly resemble a "painting by Van Gogh," demonstrating how the generated visuals align more significantly with the given textual description as the guidance scale increase to 1.
@@ 4.2 Diffusion image Inversion and Editing
Figure 9: Real image DDIM inversion results using SDv1.5. (a) Reconstructed samples after inversion by CFG and CFG++. (b) Quantitative comparison between CFG and CFG++.
We further explore the effect of CFG++ on DDIM inversion (dhariwal2021diffusion), where the source image is reverted to a latent vector that can reproduce the original image through generative process. DDIM inversion is well-known to break down in the usual CFG setting as CFG magnifies the accumulated error of each inversion step (mokady2023null), violating local linearization assumptions (wallace2023edict). Remarkably, CFG++ may mitigate this issue. This improves inversion and enables better image editing compared to inversion with the original CFG. We evaluate our method following the experimental setups in park2024energy and kim2024dreamsampler. More details follow.
Diffusion Image Inversion. Using the matched set of scales for $\omega$ and $\lambda$, we demonstrate the effect of CFG++ on the diffusion image inversion task. Specifically, we reconstruct the images after inversion and evaluate it through PSNR and RMSE, following the methodology of wallace2023edict. In Fig. 9, we illustrate the reconstructed examples (9a) from a real image from COCO data set and computed metrics (9b) for 5k COCO-2014 (lin2014microsoft) validation set. Both qualitative and quantitative evaluation demonstrate a consistent improvement on reconstruction performance induced by CFG++. Notably, DDIM inversion with CFG++ leads to consistent reconstruction of the source image across all guidance scales, while DDIM inversion with CFG fails to reconstruct it, as shown in Fig. 10 for another real image.
Image Editing. Fig. 11 compares the image editing result using CFG and CFG++ followed by image inversion. In the image editing stage, a word in the source text on the left side is swapped with the target concept, and this modified text is used as condition for sampling. For both synthetic and real images, CFG++ successfully edits the target concept while preserving other concepts, such as background. In particular, the water splashing or brush in the background, which tend to disappear in the conventional CFG, are maintained through the inversion process by CFG++. This emphasizes CFG++’s superior ability to retain specific scene elements that are frequently lost in the standard CFG approach.
Figure 11: Comparison on image editing via DDIM inversion using SDXL: (top) synthetic image editing, (bottom) real image editing.
@@ 4.3 Text-Conditioned Inverse Problems
Inverse problem involves restoring the original data ${\boldsymbol{x}}$ from a noisy measurement ${\boldsymbol{y}}={\mathcal{A}}({\boldsymbol{x}})+{\boldsymbol{n}}$, where ${\mathcal{A}}$ represents an imaging operator introduces distortions (e.g, Gaussian blur) and ${\boldsymbol{n}}$ denotes the measurement noise. Diffusion inverse solvers (DIS) address this challenge by leveraging pre-trained diffusion models as implicit priors and performing posterior sampling ${\boldsymbol{x}}\sim p({\boldsymbol{x}}|{\boldsymbol{y}})$. Methods that leverage latent diffusion have gained recent interest (rout2024solving; song2024solving), but leveraging texts for solving these problems remains relatively underexplored (chung2023prompt; kim2023regularization). One of the main reasons for this is that naively using CFG on top of latent DIS leads to diverging samples (chung2023prompt). Several heuristic modifications such as null prompt optimization (kim2023regularization) with modified sampling schemes were needed to mitigate this drawback. This naturally leads to the question: Is it possible to leverage CFG guidance as a plug-and-play component of existing solvers? Here, we answer this question with a positive by showing that CFG++ enables the incorporation of text prompts into a standard solver. Specifically, we focus on comparing the performance of PSLD (rout2024solving) combined with CFG and CFG++ in solving linear inverse problems. This evaluation utilizes the FFHQ (karras2019style) 512x512 dataset and the text prompt "a high-quality photo of a face". Further details about the experimental settings can be found in the Appendix A.
As shown in Tab. 2, our method mostly outperforms both the vanilla PSLD and PSLD with CFG. The superiority is also evident from the Fig. 12, where CFG++ consistently delivers high-quality reconstructions across all tasks. PSLD with CFG often suffer from artifacts and blurriness. Conversely, CFG++ achieves better fidelity, clearly distinguishing between faces and faithfully reproducing fine details like eyelids and hair texture. For more results, please refer to the Appendix B.
Figure 12: Qualitative comparison on various inverse problems using PSLD (rout2024solving) under CFG and CFG++.
@@
5 Related Works and Discussion
Related works. kynkaanniemi2024applying observed that applying CFG guidance at the earlier stages of sampling always led to detrimental effects and drastically reduced the diversity of the samples. The guidance at the later stages of sampling had minimal effects. Drawing upon these observations, it was empirically shown that applying CFG only in the limited time interval near the middle led to the best performance. Similar observations were made in the SD community (howard2022; birch2023) in adjusting the guidance scale across $t$. These findings are orthogonal to ours in that these methods keep the sampling trajectory the same and try to empirically adjust the strength of the guidance, while we aim to design a different trajectory.
Implications. In several studies, attempts at designing a better trajectory of the generative path were made (liu2023flow; lipman2023flow). As diffusion models can only estimate the tangent vector of the trajectory at time $t$, it is now a well regarded consensus that minimizing the curvature of the trajectory leads to faster sampling (liu2023flow; lee2023minimizing) and better distillation results (liu2023instaflow). In this regard, the excessively high curvature and oscillating behavior of CFG with high guidance scales would hamper the performance of fast sampling and distillation, which was constantly reported in the literature as difficulties in extending distillation methods (salimans2022progressive; song2023consistency) to the latent space (meng2023distillation; luo2023latent; chen2024pixart). Straightening the diffusion path prior to distillation with Reflow (liu2023flow) exhibited much better results than aiming for direct distillation. In this regard, CFG++ can also be considered a better alternative to CFG for enhanced distillation performance.

### v2

@@ 3.2 Geometry of CFG++
Table 1: Quantitative evaluation of 50NFE DDIM T2I with SD v1.5 on COCO 10k
@@ 4 Experimental Results
In this section, we design experiments to show the limitations of CFG and how CFG++ can effectively mitigate these downsides. The main experiments were conducted with SD v1.5 or SDXL with 50 NFE DDIM sampling. In this regime, we searched for the matching guidance values of $\omega$ and $\lambda$ for a fair comparison. We fix $\lambda=0.2,0.4,0.6,0.8,1.0$ and find the $\omega$ values that produce the images that are of closest proximity in terms of LPIPS distance given the same seed. We found that the corresponding values were $\omega=2.0,5.0,7.5,9.0,12.5$, respectively. Some of the experiments were also conduced with distilled model such as SDXL-turbo, lightning.
@@ 4.1 Text-to-Image Generation
Using the corresponding scales for $\omega$ and $\lambda$, we directly compare the performance of the T2I task using SD v1.5 and SDXL. In Tab. 1, we report quantitative metrics using 10k images generated from COCO captions (lin2014microsoft). Here, we observe a constant improvement of the FID metric across all guidance scales (also see Fig. 13 for an apples-to-apples comparison), with approximately the same level of CLIP similarity or better.
Model Metric CFG CFG++ (ours)
[…]
Table 2: Quant. eval. on accelerated T2I sampling
The improvements can also be clearly seen in Fig. 10 (SD v1.5), where the unnatural components of the generated images are corrected. Specifically, we see that unnatural depictions of human hands, and incorrect renderings of the text are corrected in CFG++, a long-standing research question in and of its own (podell2023sdxl; pelykh2024giving; chen2023textdiffuser). We find that the improvement gain from CFG++ is even more dramatic for distilled diffusion models such as SDXL-{turbo, lightning}. We quantify the results on the same 5k prompts with 6 NFE sampling. In Fig. 5 and Fig. 11, we see significant boosts in the quality of the generated images, which is also depicted in the improvements seen in Tab. 2.
To show the compatibility of CFG++ with higher-order solvers, we experiment with DPM++2M (lu2022dpmpp) solver with 20 NFE. We note that in such low NFE regime, a CFG scale of 5.0 corresponds to the CFG++ scale of 1.0, and we have to slightly extrapolate the value of $\lambda\geq 1.0$ to achieve stronger guidance results. In Tab. 2, we again see that CFG++ yields results that are favorable to standard CFG.
Figure 5: T2I using SDXL-Lightning, 6 NFE, CFG vs CFG++.
@@ 4.2 Diffusion image Inversion and Editing
Figure 6: Inversion and editing results. (a) Reconstructed samples after inversion by CFG and CFG++. (b) Quantitative comparison between CFG and CFG++ for reconstruction. (c) Image editing comparison via SDXL.
We further explore the effect of CFG++ on DDIM inversion (dhariwal2021diffusion), where the source image is reverted to a latent vector that can reproduce the original image through generative process. DDIM inversion is well-known to break down in the usual CFG setting as CFG magnifies the accumulated error of each inversion step (mokady2023null), violating local linearization assumptions (wallace2023edict). We show that CFG++ mitigates this issue by improving the inversion and image editing capabilities. We evaluate our method following the experimental setups in park2024energy and kim2024dreamsampler.
Diffusion Image Inversion. Using the matched set of scales for $\omega$ and $\lambda$, we demonstrate the effect of CFG++ on the diffusion image inversion task. Specifically, we reconstruct the images after inversion and evaluate it through PSNR and RMSE, following the methodology of wallace2023edict. In Fig. 6, we illustrate the reconstructed examples (6a) from a real image from COCO data set and computed metrics (6b) for 5k COCO-2014 (lin2014microsoft) validation set. Both qualitative and quantitative evaluations demonstrate a consistent improvement on reconstruction performance induced by CFG++. Notably, DDIM inversion with CFG++ leads to a consistent reconstruction of the source image across all guidance scales, while DDIM inversion with CFG fails to reconstruct it, as shown in Fig. 14 for another real image.
Image Editing. Fig. 6c compares the image editing result using CFG and CFG++ followed by image inversion with 5 NFEs. In the image editing stage, a word in the source text highlighted by green color is swapped with the target concept, and this modified text is used as the condition for sampling. CFG++ successfully edits the target concept while preserving other concepts, such as background. In particular, the water splashing in the background, which tends to disappear in the conventional CFG, is maintained through the inversion process by CFG++. This emphasizes CFG++’s superior ability to retain specific scene elements that are frequently lost in the standard CFG approach. Moreover, standard CFG’s disrupted DDIM inversion leads to saturated and less faithful editing results, whereas CFG++ delivers precise and high-fidelity edits.
@@ 4.3 Text-Conditioned Inverse Problems
Inverse problem involves restoring the original data ${\boldsymbol{x}}$ from a noisy measurement ${\boldsymbol{y}}={\mathcal{A}}({\boldsymbol{x}})+{\boldsymbol{n}}$, where ${\mathcal{A}}$ represents an imaging operator introduces distortions (e.g, Gaussian blur) and ${\boldsymbol{n}}$ denotes the measurement noise. Diffusion inverse solvers (DIS) address this challenge by leveraging pre-trained diffusion models as implicit priors and performing posterior sampling ${\boldsymbol{x}}\sim p({\boldsymbol{x}}|{\boldsymbol{y}})$. Methods that leverage latent diffusion have gained recent interest (rout2024solving; song2024solving), but leveraging texts for solving these problems remains relatively underexplored (chung2023prompt; kim2023regularization). One of the main reasons for this is that naively using CFG on top of latent DIS leads to diverging samples (chung2023prompt). Several heuristic modifications such as null prompt optimization (kim2023regularization) with modified sampling schemes were needed to mitigate this drawback. This naturally leads to the question: Is it possible to leverage CFG guidance as a plug-and-play component of existing solvers? Here, we answer this question with a positive by showing that CFG++ enables the incorporation of text prompts into a standard solver. Specifically, we focus on comparing the performance of PSLD (rout2024solving) combined with CFG and CFG++ in solving linear inverse problems. This evaluation utilizes the FFHQ (karras2019style) 512x512 dataset and the text prompt "a high-quality photo of a face". Further details about the experimental settings can be found in the Appendix C.
As shown in Tab. 3, our method mostly outperforms both the vanilla PSLD and PSLD with CFG. The superiority is also evident from the Fig. 7, where CFG++ consistently delivers high-quality reconstructions across all tasks. PSLD with CFG often suffer from artifacts and blurriness. Conversely, CFG++ achieves better fidelity, clearly distinguishing between faces and faithfully reproducing fine details like eyelids and hair texture. For more results, please refer to the Appendix D.
Figure 7: Qualitative comparison on various inverse problems using PSLD (rout2024solving) under CFG and CFG++. For more results, please refer to the Appendix D.
@@
5 Related Works and Discussions
kynkaanniemi2024applying observed that applying CFG guidance at the earlier stages of sampling always led to detrimental effects and drastically reduced the diversity of the samples. The guidance at the later stages of sampling had minimal effects. Drawing upon these observations, it was empirically shown that applying CFG only in the limited time interval near the middle led to the best performance. Similar observations were made in the SD community (howard2022; birch2023) in adjusting the guidance scale across $t$. These findings are orthogonal to ours in that these methods keep the sampling trajectory the same and try to empirically adjust the strength of the guidance, while we aim to design a different trajectory. More recently, bradley2024classifier observed that the CFG score is not a valid denoising direction, showing that in the asymptotic limit, CFG sampling can be considered as a specific type of predictor-corrector (PC) sampling, dubbed PCG. PCG shares a similar spirit with CFG++ in that for the mainstream of sampling, one defers from the use of mixed CFG score (we use unconditional score, PCG uses conditional score) to avoid invalid directions. While PCG derives an interesting connection of PC sampling with CFG, the proposed algorithm does not improve the sampling performance. In contrast, CFG++ improves the sample quality without additional computation overhead. Since CFG++ attempts to make adjustments of CFG using a linear combination of score functions, it is not surprising to see that CFG++ sampling can be achieved by setting a time varying schedule. Specifically, by setting a time-varying schedule $\omega_{t}$ as $\omega_{t}=-\gamma_{t}/\xi_{t}$, where $\gamma_{t}=\frac{\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}},\xi_{t}=\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}$, one can achieve the same effect with CFG as CFG++. However, this specific choice of time-varying guidance scale has not been reported in the literature, as it is a schedule that is relatively complex to be drawn heuristically. We further show in Appendix B that such choice removes undesirable oscillatory behavior in the evolution of the posterior mean.